Sue received 20 ore than twice as many votes as Michael. 3,794 votes were cast. How many did each candidate get?

Answer by booklover15(10)

Number of sue's votes = s
Number of michael's votes = m
s+m = 3,794
s = 2m+20
So we substitute the s
2m+20+m = 3,794
3m = 3.794 - 20
3m = 3.774
m = 1,258
So we can find s now
s+m = 3,794
s+1,258 = 3,794
s= 2,536